# DexGANGrasp: Dexterous Generative Adversarial Grasping Synthesis for Task-Oriented Manipulation

### v1

@@ V-A Experimental Setup
The experimental setup in the simulation consists of a Franka robot, Robotiq-3F hand, and a RealSense D415 camera. In the real-world experiment, the robot arm is replaced by a Kuka iiwa, as our model is robot-independent.
We segment the point cloud from the RealSense camera using RANSAC plane removal to obtain the segmented object point cloud. This point cloud is encoded using Basis Point Set (BPS) [43] and fed into DexGenerator to generate grasp candidates. The top-ranked grasp is then executed by the robot.
@@
In addition to real-world evaluation, we use data generated in simulation with Gazebo [44], following the setup in [26] for both training and evaluation. Specifically, a heuristic grasp planner is used to generate grasp data. The object models are sourced from the BIGBIRD [45], KIT [46], and YCB [47] datasets. For training in simulation, we use 115 graspable objects filtered from approximately 280 objects in the BIGBIRD and KIT datasets based on their graspability and object type. A total of 43,279 unique grasps have been executed in the simulation. With data augmentation of spawning objects in different poses, we eventually obtain 2.1 million grasps across 5750 various grasping scenes.
@@
Objects Methods Bath Detergent Curry Fizzy Tablets Instant Sauce Nut Candy Potatoe Dumpling Spray flask Tomato Soup Toy Airplane Bleach Cleanser Pitcher Base Mug Cracker Box Mustard Bottle Pudding Box Mini Soccer Ball Average Succ Rate FFHNet w/o eval $95.0\%$ $70.0\%$ $85.0\%$ $95.0\%$ $95.0\%$ $95.0\%$ $95.0\%$ $90.0\%$ $25.0\%$ $80.0\%$ $95.0\%$ $95.0\%$ $80.0\%$ $65.0\%$ $85.0\%$ $100.0\%$ $84.06\%$ DexGanGrasp w/o eval $95.0\%$ $80.0\%$ $80.0\%$ $90.0\%$ $100.0\%$ $100.0\%$ $100.0\%$ $100.0\%$ $50.0\%$ $100.0\%$ $95.0\%$ $100.0\%$ $100.0\%$ $95.0\%$ $80.0\%$ $95.0\%$ $91.25\%$ FFHNet with eval $100.0\%$ $90.0\%$ $100.0\%$ $100.0\%$ $95.0\%$ $100.0\%$ $100.0\%$ $100.0\%$ $15.0\%$ $100.0\%$ $90.0\%$ $80.0\%$ $70.0\%$ $70.0\%$ $100.0\%$ $95.0\%$ $87.81\%$ DexGanGrasp with disc $100.0\%$ $70.0\%$ $75.0\%$ $100.0\%$ $100.0\%$ $85.0\%$ $100.0\%$ $90.0\%$ $40.0\%$ $100.0\%$ $100.0\%$ $90.0\%$ $100.0\%$ $95.0\%$ $100.0\%$ $100.0\%$ $90.31\%$ DexGanGrasp with eval $100.0\%$ $100.0\%$ $85.0\%$ $100.0\%$ $100.0\%$ $100.0\%$ $100.0\%$ $100.0\%$ $45.0\%$ $95.0\%$ $100.0\%$ $100.0\%$ $100.0\%$ $100.0\%$ $95.0\%$ $100.0\%$ $95.0\%$
@@
Objects Methods Realsense Box Mustard Bottle Agile Bottle Spam Can YCB Drill Red Mug Green Pan Average Succ Rate FFHNet $90.0\%$ $70.0\%$ $90.0\%$ $80.0\%$ $60.0\%$ $70.0\%$ $50.0\%$ $72.86\%$ DexGanGrasp $100.0\%$ $100.0\%$ $90.0\%$ $90.0\%$ $80.0\%$ $100.0\%$ $80.0\%$ $91.43\%$
@@
Objects (affordance) Hammer (hand over) Hammer (use) Green Pan (grab) Spray (use) Red Brush (grab) Average Succ Rate Grasping $4/5$ $2/5$ $3/5$ $4/5$ $/5$ 5 $72.0\%$ VLMs $/6$ 6 $9/11$ $6/9$ $6/7$ $5/9$ $77.95\%$
@@ V-B Evaluation Metrics
$cos(\mathbf{\alpha})$ $=0.5\cdot trace(\mathbf{R_{ref}})-1$
@@
$Cov(\mathbf{G}_{gen},\mathbf{G}_{gt})=\frac{|\{\arg\min_{\mathbf{G}_{gt}}d(\mathbf{g}_{gen},\mathbf{g}_{gt})|\mathbf{g}_{gen}\in\mathbf{G}_{gen}\}|}{|\mathbf{G}_{gt}|}$ (11)
@@ V-C Simulation Grasping Results
We conduct grasping experiments in simulation for 16 unknown objects, 8 KIT objects, and 8 YCB objects, in table I. Each object is grasped 20 times, resulting in a total of 320 grasp trials.
Firstly, we evaluate the grasp quality alone from generative models as GANs in DexGanGrasp and Variational Autoencoder (VAE) in FFHNet, without DexEvaluator. To facilitate a fair comparison for the grasp generator without a grasp evaluator, we evaluate the grasp samplers in simulation by executing the top 20 grasps instead of the single topmost one. Here DexGanGrasp achieves a high success rate of $91.25\%$, surpassing FFHNet’s $84.06\%$. With DexEvaluator, DexGanGrasp still maintains a $7.19\%$ higher success rate than FFHNet. We observe fewer failures from DexGanGrasp in terms of unstable grasp poses due to wrong palm translation or rotation, indicating higher grasp quality both before and after filtering from DexEvaluator.
We further compare the filtering capability of grasps with either DexEvaluator or DexDiscriminator. DexEvaluator outperforms DexDiscriminator by $4.69\%$ while filtering DexGanGrasp. This difference can be attributed to the fact that DexDiscriminator is trained with only positive grasps and generated grasps from DexGenerator, while DexEvaluator is trained on both positive and negative grasps from the ground truth.
@@ V-D Real world Grasping Results
Following the simulation experiments, we conduct the real-world grasping experiment. We test 7 unknown objects with DexGanGrasp and FFHNet, both paired with DexEvaluator. Each object is grasped 10 times resulting in a total of 70 grasp trials for each model in table II. In the real-world experiment, DexGanGrasp achieves the best success rate of $91.43\%$ outperforming FFHNet by $18.57\%$. For real-time performance, DexGanGrasp achieves an inference speed of $30ms$, tested in a commercial laptop with RTX 4090 16GB GDDR6 GPU.
@@ V-F Real World Task-Oriented Grasping Results
For the task-oriented grasping with DexAfford-Prompt, we choose 4 tool-related objects and test DexAfford-Prompt  with 3 different prompts ‘hand over’, ‘use’ and ‘grab’. Each object with a different prompt is tested 5 times, resulting in a total of 25 grasps, with details in table III
DexAfford-Prompt achieves an overall grasping success rate of $72.0\%$. For the hammer, DexAfford-Prompt enables two different prompts: ‘hand over’, where the robot is supposed to grasp the top part with the predicted part name of ‘hammer head’ and leave the handle for another person or robot to take over; and ‘use’, where the robot is supposed to grasp the bottom handle with the predicted part name of ‘hammer handle’.
Besides grasping, VLMs for part segmentation achieves an average success rate of $77.95\%$ in table III. The number of attempts for VLMs are often more than 5 times because certain grasps are not executed successfully for example the robot goes into singularity or reaches joint limit. We observed several failures from VLMs, which can occasionally predict the wrong part, detailed in figure 4. Reasons could be the different lighting conditions compared the training data in VLMs.
Figure 4: Successful part predictions on first row and some failure cases in second row from VLMs VLPart [42].
@@ V-G Ablation Study
We conduct an ablation study to explore various design choices and loss functions for our DexGanGrasp, as well as comparing it with the baseline FFHNet in table IV. We formulate different architectures and training hyperparameters and apply the proposed MAGD and Cov metric cumulatively over a batch of 64 randomly selected items and their corresponding grasp distributions from our evaluation dataset.
The initial implementation is a vanilla GAN without distance losses $L_{dist}$. This implementation performs poorly due to the absence of guidance provided by the distance losses. Next, we test a GAN that includes distance losses but increases the Discriminator’s training rate, omitting the label-switching approach mentioned in section IV-A. This configuration yields slightly better results than the vanilla GAN but still lacks diversity as indicated by the Cov metric. Subsequently, we test a Wasserstein GAN (WGAN) [49] that utilizes the Wasserstein distance as a loss function, measuring the minimum work required to transform one probability distribution into another. This approach is highly unstable and only shows improvement in the Cov metric. To stabilize training, a gradient penalty was applied [50], improving results in coverage, rotation, and joint configuration, although there was still a significant error in translation. WGAN-GP improves upon the original WGAN by incorporating a gradient penalty to address gradient-related issues associated with weight clipping.
The final variant implements a GAN with distance losses and reduces training frequency for DexDiscriminator on generated grasps. With this approach, we identify two model checkpoints that outperform FFHNet in different categories. Epoch 32 excelled in all metrics except Cov, while Epoch 37 achieved the best results in translation and coverage overall. For our experiments, we select Epoch 32 due to its superior performance in crucial metrics for a good grasp (translation, rotation and joint configuration). Simplistically, coverage measures overlap with the ground truth distribution, which does not necessarily indicate a better distribution.
@@
Cumulative MAGD(w/o eval) Ablated Models Translation $\downarrow$ Rotation $\downarrow$ Joints $\downarrow$ Cov $\uparrow$ FFHNet 30.22 23779.96 0.8621 29.47% DexGanGrasp (ep = 37) 25.75 25587.59 1.6773 29.77% DexGanGrasp (ep = 32) 26.08 22656.06 0.7249 25.91% GAN (no label switch) 46.52 37683.35 5.5801 0.88% Wasserstein Gan with GP 58.65 26496.25 0.9936 18.07% Wasserstein Gan 71.10 38737.16 109.4223 1.04% Vanilla GAN (without reconstruction losses) 166.49 38879.15 320.90 0.47%
@@ VI Conclusion
In this work, we propose DexGanGrasp for the dexterous grasping problem utilizing conditional GANs to grasp unknown objects with a single view in real time. The proposed DexGenerator  in DexGanGrasp demonstrates its ability to generate higher quality and more diverse grasps compared to baseline FFHNet, assessed through our proposed MAGD and Cov metrics and extensive simulation as well as real-world evaluations. To address task-oriented grasping, we further extend DexGanGrasp  to DexAfford-Prompt, enabling an open-vocabulary affordance grounding pipeline with successful real-world deployments to unknown objects, showcasing the potential for future advancements in task-based manipulation.
We have identified several limitations in the current pipeline. Firstly, when dealing with slightly transparent objects such as sprays, the point cloud from the RealSense camera can be very noisy on the surface, leading to an unrealistic grasp distribution from DexGanGrasp. Secondly, the current state-of-the-art VLMs needs to be more stable for affordance prediction. An affordance-aware grasp generation model is a promising future direction.

### v2

@@ V-A Experimental Setup
The experimental setup in the simulation consists of a Franka robot, Robotiq-3F hand, and a RealSense D415 camera. In the real-world experiment, the robot arm is replaced by a Kuka iiwa, as our model is robot-independent. The robot grasping workspace is calibrated via homography [47].
We segment the point cloud from the RealSense camera using RANSAC plane removal to obtain the segmented object point cloud. This point cloud is encoded using Basis Point Set (BPS) [48] and fed into DexGenerator to generate grasp candidates. The top-ranked grasp by DexEvaluator is then executed by the robot.
@@
We generate a synthetic grasping dataset for Robotiq-3F hand in simulation Gazebo [49], following the setup in [26] for both training and evaluation. Specifically, a heuristic grasp planner is used to generate grasp data. The object models are sourced from the BIGBIRD [50], KIT [51], and YCB [52] datasets. For training in simulation, we use 115 graspable objects filtered from approximately 280 objects in the BIGBIRD and KIT datasets based on their graspability and object type. A total of 43,279 unique grasps have been executed in the simulation. With data augmentation of spawning objects in different poses, we eventually obtain 2.1 million grasps across 5750 various grasping scenes.
@@
Objects Methods Bath Detergent Curry Fizzy Tablets Instant Sauce Nut Candy Potatoe Dumpling Spray flask Tomato Soup Toy Airplane Bleach Cleanser Pitcher Base Mug Cracker Box Mustard Bottle Pudding Box Mini Soccer Ball Average Succ Rate FFHNet w/o eval $95.0\%$ $70.0\%$ $85.0\%$ $95.0\%$ $95.0\%$ $95.0\%$ $95.0\%$ $90.0\%$ $25.0\%$ $80.0\%$ $95.0\%$ $95.0\%$ $80.0\%$ $65.0\%$ $85.0\%$ $100.0\%$ $84.06\%$ DexGanGrasp w/o eval $95.0\%$ $80.0\%$ $80.0\%$ $90.0\%$ $100.0\%$ $100.0\%$ $100.0\%$ $100.0\%$ $50.0\%$ $100.0\%$ $95.0\%$ $100.0\%$ $100.0\%$ $95.0\%$ $80.0\%$ $95.0\%$ $91.25\%$ FFHNet with eval $100.0\%$ $90.0\%$ $100.0\%$ $100.0\%$ $95.0\%$ $100.0\%$ $100.0\%$ $100.0\%$ $15.0\%$ $100.0\%$ $90.0\%$ $80.0\%$ $70.0\%$ $70.0\%$ $100.0\%$ $95.0\%$ $87.81\%$ DexGanGrasp with disc $100.0\%$ $70.0\%$ $75.0\%$ $100.0\%$ $100.0\%$ $85.0\%$ $100.0\%$ $90.0\%$ $40.0\%$ $100.0\%$ $100.0\%$ $90.0\%$ $100.0\%$ $95.0\%$ $100.0\%$ $100.0\%$ $90.31\%$ DexGanGrasp with eval $100.0\%$ $100.0\%$ $85.0\%$ $100.0\%$ $100.0\%$ $100.0\%$ $100.0\%$ $100.0\%$ $45.0\%$ $95.0\%$ $100.0\%$ $100.0\%$ $100.0\%$ $100.0\%$ $95.0\%$ $100.0\%$ $\mathbb{95.0\%}$
@@
Objects Methods Realsense Box Mustard Bottle Agile Bottle Spam Can YCB Drill Red Mug Green Pan Average Succ Rate Inference Speed FFHNet $90.0\%$ $70.0\%$ $90.0\%$ $80.0\%$ $60.0\%$ $70.0\%$ $50.0\%$ $72.86\%$ $30ms$ DexGanGrasp $100.0\%$ $100.0\%$ $90.0\%$ $90.0\%$ $80.0\%$ $100.0\%$ $80.0\%$ $\mathbb{91.43\%}$ $30ms$
@@
Objects (affordance) Hammer (hand over) Hammer (use) Green Pan (grab) Spray (use) Red Brush (grab) Average Succ Rate Grasping $4/5$ $2/5$ $3/5$ $4/5$ $/5$ 5 $72.0\%$ VLM [44] $/6$ 6 $9/11$ $6/9$ $6/7$ $5/9$ $77.95\%$
@@ V-B Evaluation Metrics
$cos(\mathbf{\alpha})$ $=(trace(\mathbf{R_{ref}})-1)/2$
@@
$Cov=\frac{|\{\arg\min_{\mathbf{G}_{gt}}d(\mathbf{g}_{gen},\mathbf{g}_{gt})|\mathbf{g}_{gen}\in\mathbf{G}_{gen}\}|}{|\mathbf{G}_{gt}|}$ (11)
@@ V-C Simulation Grasping Results
We conduct grasping experiments in simulation for 16 unknown objects, 8 KIT objects, and 8 YCB objects in table I. Each object is grasped 20 times, resulting in a total of 320 grasp trials.
Firstly, we evaluate the grasp quality alone from generative models as GAN in DexGanGrasp and Variational Autoencoder (VAE) in FFHNet, without DexEvaluator. To facilitate a fair comparison for the grasp generator without a grasp evaluator, we evaluate the grasp samplers in simulation by executing the top 20 grasps instead of the single topmost one. Here DexGanGrasp achieves a high success rate of $91.25\%$, surpassing FFHNet’s $84.06\%$. With DexEvaluator, DexGanGrasp still maintains a $7.19\%$ higher success rate than FFHNet. We observe fewer failures from DexGanGrasp in terms of unstable grasp poses due to wrong palm translation or rotation, indicating higher grasp quality both before and after filtering from DexEvaluator.
We further compare the filtering capability of grasps with either DexEvaluator or DexDiscriminator. DexEvaluator outperforms DexDiscriminator by $4.69\%$ while filtering DexGanGrasp. This difference can be attributed to the fact that DexDiscriminator is evaluating grasps based on realism (trained on $\mathbf{g}_{pos}$ and $\mathbf{g_{gen}}$), while DexEvaluator is evaluating grasps based on probability of grasping success (trained on $\mathbf{g}_{pos}$ and $\mathbf{g}_{neg}$). This means the discriminator itself cannot be directly used as the evaluator as $\mathbf{g_{gen}}$ can be either positive or negative grasps.
@@ V-D Real world Grasping Results
Following the simulation experiments, we conduct the real-world grasping experiment. We test 7 unknown objects with DexGanGrasp and FFHNet, both paired with DexEvaluator. Each object is grasped 10 times, resulting in a total of 70 grasp trials for each model in table II. In the real-world experiment, DexGanGrasp achieves the best success rate of $91.43\%$ outperforming FFHNet by $18.57\%$. For real-time performance, DexGanGrasp achieves an inference speed of $30ms$, tested in a commercial laptop with RTX 4090 16GB GDDR6 GPU.
@@ V-F Real World Task-Oriented Grasping Results
For the task-oriented grasping with DexAfford-Prompt, we choose 4 tool-related objects and test them with 3 different prompts ‘hand over’, ‘use’ and ‘grab’. Each object with a different prompt is tested 5 times, resulting in a total of 25 grasps, with details in table III
DexAfford-Prompt achieves an overall grasping success rate of $72.0\%$. For the hammer, it enables two different prompts: ‘hand over’, where the robot is supposed to grasp the top part with the predicted part name of ‘hammer head’ and leave the handle for another person or robot to take over; and ‘use’, where the robot is supposed to grasp the bottom handle with the predicted part name of ‘hammer handle’.
Besides grasping, the VLM achieves an average success rate of $77.95\%$ for part segmentation in table III. The number of attempts for the VLM are often more than 5 times because certain grasps are not executed successfully. For example, the robot goes into a singularity or reaches the joint limit. We observed several failures from the VLM, which can occasionally predict the wrong part, detailed in figure 4. The reasons could be the different lighting conditions compared to the training data in the VLM.
Figure 4: Successful part predictions on the first row and some failure cases in the second row from VLPart (VLM[44].
@@ V-G Ablation Study
We perform an ablation study to assess various design choices and loss functions for our DexGanGraspcompared to the baseline FFHNet (table IV). We test different architectures and hyperparameters, applying the MAGD and Cov metrics over a random batch of 64 items and their grasp distributions.
The final DexGanGrasp  variant employs a cGAN with distance losses and reduced discriminator training (section IV-E). We identify two optimal checkpoints: Epoch 32, excelling in all metrics except Cov, and Epoch 37, which maximizes translation and coverage. We choose Epoch 32 because of its superior performance in key grasp metrics (translation, rotation, joint configuration).
We also test a vanilla cGAN without distance losses $\mathcal{L}{dist}$, which performs poorly. Adding $\mathcal{L}{dist}$ and increasing discriminator updates (omitting label-switching) improves results but lacks diversity in Cov. Next, we evaluate a Conditional Wasserstein GAN (cWGAN)[54] using the Wasserstein distance as a loss. This approach is unstable, improving only Cov. Stabilizing with gradient penalties (cWGAN-GP[54]) yields better results in coverage, rotation, and joint configuration, though translation errors persist.
@@
Cumulative MAGD(w/o eval) Ablated Models Translation $\downarrow$ Rotation $\downarrow$ Joints $\downarrow$ Cov $\uparrow$ FFHNet 30.22 23779.96 0.8621 29.47% DexGanGrasp (ep = 37) 25.75 25587.59 1.6773 29.77% DexGanGrasp (ep = 32) 26.08 22656.06 0.7249 25.91% Wasserstein cGAN with GP 58.65 26496.25 0.9936 18.07% Wasserstein cGAN 71.10 38737.16 109.4223 1.04% cGAN (increased train rate of discriminator) 46.52 37683.35 5.5801 0.88% Vanilla cGAN (without distance losses) 166.49 38879.15 320.90 0.47%
@@ VI Conclusion
In this work, we propose DexGanGrasp for the dexterous grasping problem utilizing cGAN to grasp unknown objects with a single view in real-time. The proposed DexGenerator in DexGanGrasp demonstrates its ability to generate higher quality and more diverse grasps compared to baseline FFHNet, assessed through our proposed MAGD and Cov metrics and extensive simulation as well as real-world evaluations. To address task-oriented grasping, we further extend DexGanGrasp to DexAfford-Prompt, enabling an open-vocabulary affordance grounding pipeline with successful real-world deployments to unknown objects, showcasing the potential for future advancements in task-based manipulation.
We have identified several limitations in the current pipeline. Firstly, when dealing with slightly transparent objects such as sprays, the point cloud from the RealSense camera can be very noisy on the surface, leading to an unrealistic grasp distribution from DexGanGrasp. Secondly, the current state-of-the-art VLM needs to be more stable for affordance prediction. An affordance-aware grasp generation model is a promising future direction.